### v2

# Learning Unified User Quantized Tokenizers for User Representation

Abstract.
Multi-source user representation learning plays a critical role in enabling personalized services on web platforms (e.g., Alipay). While prior works have adopted late-fusion strategies to combine heterogeneous data sources, they suffer from three key limitations: lack of unified representation frameworks, scalability and storage issues in data compression, and inflexible cross-task generalization. To address these challenges, we propose U²QT (Unified User Quantized Tokenizers), a novel framework that integrates cross-domain knowledge transfer with early fusion of heterogeneous domains. Our framework employs a two-stage architecture: first, we use the Qwen3 Embedding model to derive a compact yet expressive feature representation; second, a multi-view RQ-VAE discretizes causal embeddings into compact tokens through shared and source-specific codebooks, enabling efficient storage while maintaining semantic coherence. Experimental results showcase U²QT’s advantages across diverse downstream tasks, outperforming task-specific baselines in future behavior prediction and recommendation tasks while achieving efficiency gains in storage and computation. The unified tokenization framework enables seamless integration with language models and supports industrial-scale applications.
1. Introduction
In the web platforms such as Alipay111https://global.alipay.com/platform/site/ihome, users generate diverse multi-source information, including textual profiles, behavioral histories, and structured tabular attributes. Effectively leveraging this heterogeneous data is crucial for advancing personalized services and optimizing user experience. To this end, prior studies have predominantly adopted a late-fusion strategy, wherein source-specific features are independently extracted and subsequently combined at a later stage (Purificato et al., 2024; Wu et al., 2019; Tang et al., 2021). For example, early approaches, such as user2Vec (Hallac et al., 2019), adapted document embedding models (e.g., doc2vec) to encode social media users, typically by aggregating user vectors within the same category to form category-level representations. To address the limitation of simple aggregation, subsequent works designed more expressive fusion mechanisms. For instance, MFN (Bian et al., 2021) introduced a Macro-micro Fusion Network that captures semantic alignment at the user level using a category-based fusion component. In parallel, attention-based models like MSSR (Lin et al., 2024) explored how to jointly leverage both inter- and intra-sequence user interactions to capture a user’s diverse behavioral patterns.
With the rapid growth of large-scale pre-training, general-purpose user representation frameworks have emerged. For example, One4all (Shin et al., 2021) learns universal user embeddings via large-scale pre-training, and its extension U-MLP further enhances user targeting via an MLP-based decoder. Sequence modeling methods such as MSDP (Fu et al., 2023) focus on modeling temporal dynamics of user behaviors at multiple scales.
Recently, to support zero-shot transfer and few-shot user targeting, FOUND (Dou et al., 2025) proposed a self-supervised pre-training framework with user-text alignment, advancing user modeling towards more generalizable and transferable representations.
However, such approaches face three critical limitations:
(i) Lack of unified representation for heterogeneous data: As depicted in Figure 1, methods like User2Vec overlook cross-source interactions by relying on single-source features, while existing approaches (e.g., MSSR, MFN) employ late fusion strategies in multi-source integration, limiting cross-source interaction potential.
(ii) Data utilization and storage issues: As shown in Figure 2, methods like FOUND require massive scale to store embeddings for every user, leading to prohibitive storage demands and further impede the model’s capacity to fully capitalize on abundant information.
(iii) Inflexible cross-task generalization: Methods such as Conure (Yuan et al., 2021) and TERACON (Kim et al., 2023) require additional fine-tuning of a large number of user embeddings when adapted to new tasks, resulting in limited flexibility and restricted practical applicability.
To address the above challenges, we propose Unified User Quantized Tokenizers (U2QT), which compresses the heterogeneous multi-source data into a unified quantitative language for efficient data storage, utilization and computation. Specifically, we adopt a two-stage quantitative language-based data compression framework to solve Challenge (i) and (ii). In the first stage, heterogeneous source-specific features (e.g., text, sequential behaviors, tabular data) are projected into a unified language space in Qwen3 embedding model (Zhang et al., 2025b), which transforms these features in the format of long context into compact yet expressive embeddings. In the second stage, we propose a MRQ-VAE (Multi-View Residual Quantized Variational AutoEncoder) to discretize the causal embeddings into compact tokens stored in shared and source-specific codebooks where cross-source semantic coherence and source-unique patterns can be learned. This quantization process achieves 84× memory reduction compared to FOUND (Dou et al., 2025), with 3.5× faster training speed under equivalent training data scales, while enhancing data utilization efficiency through semantic-preserving token compression. The method enables scalable deployment of large-scale user behavior data without compromising semantic fidelity. Furthermore, due to its lightweight and quantized design, which eliminates redundant and irrelevant information, our method achieves strong cross-task adaptability. To assess the effectiveness and cross-task adaptability of our approach, we introduce a multi-task e-commerce benchmark encompassing user profile, future behavior prediction, and downstream recommendation tasks. Extensive experiments demonstrate superior adaptability and scalability across diverse tasks compared to state-of-the-art baselines. In summary, our contributions are listed as follows,
We propose unified user quantized tokenizers that integrate cross-domain knowledge transfer capabilities with heterogeneous source early fusion, enabling robust modeling of multi-source behavioral data.
Our proposed U2QT compresses the heterogeneous data into compact tokens achieving 84× memory reduction compared to SOTA models, with 3.5× faster training speed under equivalent training data scales.
Our framework allows for flexible cross-task generalization and achieves superior cross-task generalization through consistent performance gains across downstream tasks in our introduced multi-task e-commerce benchmark compared to state-of-the-art baselines.
2. Related Work
2.1. Quantitative Language
Quantitative language has been widely used in VLLM (Li et al., 2025), generative recommendation and retrieval (Li et al., 2024), which is also called semantic IDs (Zheng et al., 2024; Pan et al., 2024; Zhang et al., 2025a). For instance, VQ-Rec (Hou et al., 2023) proposes to generate code using content information for item representation. Tiger (Rajput et al., 2023) utilizes RQ-VAE to generate semantic IDs for hierarchical item representation. To represent multi-modal features, MMGRec (Liu et al., 2024) utilize graph RQ-VAE to generate semantic IDs for multi-modal item representation fused by GCN. MQL4GRec (Zhai et al., 2025) transforms item content from different domains and modalities into a unified quantitative language to facilitate the effective transfer of recommendation knowledge. SEED (Ge et al., 2023) proposes a discrete image tokenizer, designed based on the premise that visual tokens compatible with LLMs should capture high-level semantics while being generated with a 1D causal dependency.
2.2. Universal User Representation
User modeling, also known as user profiling, refers to the process of collecting, extracting, and representing user features and personal characteristics to construct comprehensive and accurate models of users(Purificato et al., 2024). Existing works has focused on the creating universal user representation(Yuan et al., 2021; Shin et al., 2021). DUPN (Ni et al., 2018) perform multitask representation learning using an attention-based RNN architecture to capture in-depth representations of portal users. PeterRec (Yuan et al., 2020) propose a parameter-efficient transfer learning framework that reduces computational costs in fine-tuning. It supports five downstream tasks, including user profile prediction and cold-start recommendations for browsers. SUMN (Gu et al., 2021) propose a Self-supervised User Modeling Network to learn universal user representations from behavior data. It consists of a self-supervised learning objective to capture and preserve meaningful user information, along with a multi-hop aggregation layer that enhances the model’s ability to integrate diverse behavioral signals. Conure (Yuan et al., 2021) propose the first continual user representation learning framework that adapts to new tasks while retaining knowledge from previous ones through iterative pruning of less important model weights. TERACON (Kim et al., 2023) propose a continual user representation learning method that captures task relationships via task-specific soft masks. A knowledge retention module with pseudo-labeling helps prevent catastrophic forgetting during sequential learning.
3. Preliminaries
3.1. Notations
In the real-world application, extensive non-sensitive user information and interactions are accessible. For instance, user behavioral sequence data is denoted as , comprising PayBill , Super Position Model , MiniProgram , and App . Here, denotes the number of users and . , , and represent the raw text for PayBill information, Super Position Model details, MiniProgram, and App descriptions, respectively, where is the token dictionary, and indicates the sequence length. Furthermore, tabular data is denoted as , where represents the number of features and denotes the dimensionality of each feature. Search text data is represented as . Each user’s information can be represented as , where and is the joint space of all possible user data. This work focuses on learning unified and robust user representations from these multi-source, multi-source data to support diverse downstream tasks.
3.2. Problem Definition
To develop a unified user representation, extensive user data can be utilized to pretrain within a unified quantitative language space:
where is the multi-source user data used for pretraining, represents the pretrained unified user representation, and denotes the proposed pretraining task, which will be illustrated in the Section 4.
With the pretrained unified user representations , we can utilize them for various downstream tasks:
where is the predicted label of -th user in the downstream task , is the model parameters for specific tasks. The specific application of our pretrained unified user representation to downstream tasks will be detailed in Section 4.4. In summary, this is a two-stage framework in which user representations are generated first and fed into a specific model for downstream task.
4. Methodology
In this section, we elaborate on the proposed Unified User Quantized Tokenizers, a novel approach of user modeling. First, we provide an overview of U2QT in Section 4.1 to facilitate a thorough understanding of our method. Subsequently, we illustrate the detail architecture of our proposed method including Multi-source Encoding (Section 4.2), and Quantized User Tokenizers (Section 4.3). Finally, we show the flexible application of user tokenizers in future behavior prediction and downstream recommendation tasks in Section 4.4. Comprehensive architecture of U2QT is illustrated in Figure 3.
4.1. Overview of Unified User Tokenizer
In this subsection, we illustrate the process of transforming complex mutli-source content of user into a concise and unified quantitative language. To better construct a unified and rich-semantic user tokenizer, we introduce a two-stage compression process, including expressive language space and discrete embedding space compression. The first stage ensures a compact yet expressive semantic representation among multi-source heterogeneous data, and the second stage contributes to efficient storage and advanced data utilization proficiency.
4.2. Multi-source Encoding
Given the heterogeneity and information disparity across data sources, prior methods typically processed each source independently and encoded it into a latent representation to streamline downstream procedures (Huang et al., 2020; Devlin et al., 2019). While this design simplifies the pipeline, it leaves a semantic gap across sources that these models struggle to bridge. To address this, we leverage LLMs to map multi-source data into a shared language space, producing richer and better-aligned embeddings. We use the Qwen3 Embedding model (Zhang et al., 2025b) to encode multi-source data. We construct a prompt that combines the query with the corresponding source data and feed it to the model. Specifically, we show an example of SPM data as illustrated in Appendix D.
Formally, final embeddings of multi-source data are derived from the hidden state of the last layer corresponding to [EOS] token as follows:
Where denotes the final representation of the multi-source input, and indexes the source. This representation compresses very long contextual information into a single embedding.
4.3. Quantized User Tokenizers
After extracting the effective information from various source data, our aim is to further compress the representation from continuous space to discrete space. Existing approaches require considerable storage and computational resources to maintain pretrained user representations. Consequently, there is a pressing need for effective data compression techniques that can significantly decrease storage and computation requirements without compromising the representational capacity of user embeddings. Motivated by the heterogeneous multi-source nature of user data, we propose a Multi-view RQ-VAE (MRQ-VAE) with a shared-specific codebook hierarchy to compress latent representations into a discrete embedding space while explicitly modeling the heterogeneous multi-source nature of user data.
For each source , the transformed source-specific user behavior is defined as , where denotes the learnable query vectors. The transformation is performed by first applying a mean pooling function for semantic summarization, followed by a shared multi-layer perceptron for dimensionality alignment. Then, MRQ-VAE first encodes it into a latent representation . The quantization then occurs in two hierarchical stages: (i) Shared Codebook Stage ( levels): Captures cross-source semantic coherence using a shared codebook , and (ii) Source-specific Codebook Stage ( levels): Refines source-unique patterns using source-specific codebooks for each source .
To reduce notational clutter, we omit symbols denoting source information and whether the codebook is shared or specific. The residual quantization process at level for user is defined as:
where is the selected code index, is a vector from the relevant codebook ( or ), and represents the residual vector. The initial residual is set to .
After processing through all levels, where the first levels are shared and the remaining levels are specific to the source , the quantized representation is obtained by summing the selected codebook vectors:
This design hierarchically disentangles common knowledge (learned via shared codebooks) from source-unique patterns (learned via specific codebooks).
The quantized representation is then fed into a source-specific MLP decoder to reconstruct the Qwen3 embedding . For notational simplicity, we also omit symbols indicating source information and whether the codebook is shared or specific. The loss function comprises:
where denotes the stop-gradient operator and is a loss coefficient. The total loss for training the user tokenizer is , ensuring both accurate reconstruction of raw content and minimization of the distance between residuals and their assigned codebook vectors.
Once the unified user tokenizer is trained, we can translate the multi-souce heterogeneous data into a unified quantitative language space. For example, if we assign a two-level shared codebooks and two-level specific codebooks for each domain, we can represent a user as , where means the token from -th level shared codebook and denotes the token from -th level -specific codebook. However, translating the user representation into a unified quantitative language may lead to a collision in which different users are assigned the same tokens. To handle with the collision, we add a special token at the end of original user tokenizers that are assigned the same tokens according to their rank in user engagement metrics:
where is the sorting function for designating a unique special token to user-defined tokenizers that generate overlapping or identical tokens, ensuring distinct identification in downstream processing. is the representation of each special token. Thus, the final user tokenizer is composed of domain-specific and special tokens, as .
4.4. Cross-Task Application
To validate the cross-task generalization capability of our proposed unified user quantized tokenizer (U2QT), we illustrate the application of U2QT in future behavior prediction and downstream recommendation task.
4.4.1. Future Behavior Prediction
In this subsection, we aim to endow our user tokenizer with the future behavior prediction capabilities. Inspired by CLIP and it’s variants (Li et al., 2022, 2023; Zhu et al., 2025), we align user tokenizer representations with future behaviors via contrastive learning for prediction. Future behavior descriptions are encoded into semantic embeddings using a LoRA-tuned (Hu et al., 2022) LLM:
where represents the process of generating user description according to the template as follow: The user purchased {items} amounting more than {num} dollars with {status} payment. In the above template, the placeholders are represented in {}: “items” denotes the purchased items, “num” represents the amount spent by the user, and “status” indicates whether the bill payment was successful. Templates for other data sources follow similar structures.
After obtaining the fused representation of user tokenizer and corresponding text embedding of future user behavior in natural langugage semantic , we employ the contrastive loss Info-NCE (van den Oord et al., 2019) to align these two representations in different semantic space. The loss is applied to pulling together the two representations from same user while pushing away two representations from different users. The contrastive loss is performed as follow:
where represents batch size and is the cosine similarity function.
4.4.2. Recommendation Task
In this section, we illustrate the usage of user tokenizer in downstream recommendation task. To describe the user-side feature, we employ the unified user tokenizer as a part of model input to representation multifaceted user information extracted from heterogeneous multi-source data as depicted in Figure 7. The overall process is as follows:
where denotes the user tokenizer of extracted from cashe storage, denotes the feature representation of and (Wang et al., 2021) is a classical method in CTR prediction.
5. Experiment
In this section, we conduct comprehensive experiments on our proposed U2QT, to demonstrate it’s effectiveness in multi-source user representation, flexible cross-task generalization and efficiency in data utilization and storage.
| Dataset | No. | Domain | Scenario | Number |
| - | General | General | 20 million | |
| 1 | future behavior | willing for takeout | 50,000 | |
| 2 | future behavior | purchasing power | 50,000 | |
| 3 | security | money Laundering | 50,000 | |
| 4 | user profile | preference for games | 50,000 |
| Methods | #1 | #2 | #3 | #4 |
| One4all | 0.7663/0.3998 | 0.8348/0.4885 | 0.8855/0.6266 | 0.9750/0.8600 |
| MSDP | 0.7891/0.4361 | 0.8464/0.4989 | 0.8810/0.6194 | 0.9731/0.8505 |
| FOUND | 0.8315/0.5144 | 0.9381/0.7228 | 0.9026/0.6952 | 0.9425/0.8064 |
| One4all+UT | 0.7752/0.4185 | 0.8649/0.5360 | 0.8961/0.6476 | 0.9658/0.8429 |
| MSDP+UT | 0.7935/0.4427 | 0.8526/0.5087 | 0.8964/0.6479 | 0.9744/0.8588 |
| U2QT | 0.8399/0.5262 | 0.9516/0.7585 | 0.9639/0.8342 | 0.9797/0.8747 |
| Data | #1 | #2 | #3 | #4 |
| All Sources | 0.8399/0.5262 | 0.9516/0.7585 | 0.9639/0.8342 | 0.9797/0.8747 |
| (-)Tabular | 0.8249/0.4988 | 0.9191/0.6756 | 0.9524/0.8045 | 0.9780/0.8675 |
| (-)App | 0.8226/0.4953 | 0.9081/0.6521 | 0.9500/0.7991 | 0.9779/0.8674 |
| (-)MiniProgram | 0.8193/0.4907 | 0.8976/0.6313 | 0.9492/0.7966 | 0.9635/0.8190 |
| (-)Search | 0.8172/0.4872 | 0.8917/0.6184 | 0.9474/0.7935 | 0.9629/0.8136 |
| (-)Bill | 0.6200/0.1697 | 0.7883/0.4316 | 0.9314/0.7701 | 0.9619/0.8135 |
5.1. Experimental Setup
5.1.1. Datasets
We pre-train our industrial framework of unified user tokenizers on a real-world dataset from Alipay that aggregates heterogeneous, multi-scenario user data. The training set comprises approximately 20,000,000 user-behavior records, each organized according to the Notations. The user-representation benchmark covers four scenarios spanning security risk control, future behavior prediction, and user profile prediction. The test set contains approximately 50,000 user-behavior records.
For the downstream recommendation task, we adopt a chronological partition: interactions from the first 21 days (approximately 1,550,000 samples) form the training set, while the subsequent 3 days (approximately 100,000 samples) are split evenly into validation and test sets using stratified random sampling. This temporal split preserves interaction order to prevent data leakage and enhances robustness through dedicated holdout periods.
5.1.2. Detailed Experimental Setup
The detailed experimental setup, including baseline descriptions, evaluation metrics, and implementation details, is provided in Appendix A, Appendix B, and Appendix C, respectively.
| Feature | Backbone | AUC | HR@10 | HR@20 |
| User Profile | Masknet | 0.6476 | 0.2308 | 0.2968 |
| (+)Tokenizer | Masknet | 0.6527 | 0.2678 | 0.4913 |
5.2. Performance Analysis
As illustrated in Table 2, the experimental results demonstrate that U2QT significantly outperforms baseline methods across downstream tasks in our designed benchmark. Our model achieves superior performance in AUC and KS in all domains and outperforms the second-best baseline 3% / 9% on average in AUC / KS. Moreover, we utilize the pretrained user tokenizer as supplementary feature in existing methods(e.g., One4all and MSDP) to verify the effectiveness. The result illustrates that these baselines gain huge improvement and outperform the original performance 1.1%/2.9% at AUC/KS for One4all and 0.7%/2.2% at AUC/KS for MSDP.
Moreover, we utilize our pretrained user tokenizer as user side feature in downstream recommendation task to verify the generalization of our proposed method. The result is illustrated in Table 4 that using tokenizer as a complementary feature achieves better performance and outperforms the baseline 0. 7% / 16. 0% / 65. 5% at AUC / HR@10 / HR@20.
Our unified user tokenizer demonstrates superior performance across a suite of benchmarks, including future behavior prediction, user profile prediction, and security analysis, as well as in downstream recommendation tasks. This success underscores its powerful ability to represent heterogeneous data and its robust generalization capability across diverse tasks.
5.3. Ablation Study
5.3.1. Data Source
In this subsection, we conduct an ablation study to evaluate the contribution of each heterogeneous data source. The experiment involves systematically removing one data domain at a time and then pre-training a corresponding tokenizer for evaluation on our benchmarks. The results, summarized in Table 3, lead to the following key conclusions:
i) All data domains contribute positively to the tokenizer’s overall effectiveness. As shown in Table 3, removing any single data source results in a performance degradation across all benchmark tasks. This underscores the synergistic value of integrating diverse data.
ii) Domain-specific data is particularly crucial for its corresponding downstream tasks. A salient example is the removal of the ”Bill” data, which led to a significant performance drop on the ”Willingness for Takeout” and ”Purchasing Power” prediction tasks. This is likely because the bill data contains rich, explicit signals about purchasing behaviors, such as takeout orders and other item expenditures, which are directly relevant to these specific predictions.
| Design | Volume | AUC | KS |
| 4share | 256128 | 0.8121 | 0.5007 |
| 2share+2specific | 256128 | 0.8205 | 0.5167 |
| 4share+2specific | 256128 | 0.8220 | 0.5204 |
| 3share+2specific | 1024256 | 0.8289 | 0.5332 |
5.3.2. Codebook Design
In this section, we present an ablation study to evaluate the effect of codebook architecture. Specifically, we compare a vanilla RQ-VAE with our proposed Multi-view RQ-VAE, and we additionally examine how performance scales with codebook capacity (number of layers, entries per layer, and code dimension). As shown in Table 5, the results yield two main observations:
i) Our Multi-view RQ-VAE consistently outperforms the vanilla RQ-VAE across all benchmarks. In the first two rows of Table 5 (which hold the configuration fixed at 256 entries per layer and 128-dimensional code vectors), the proposed architecture achieves superior results, validating the effectiveness of its multi-view design.
ii) Performance scales positively with increased codebook capacity. Comparing the 256x128 configurations, moving from ”2share+2specific” to ”3share+2specific” yields gains on both metrics, indicating that greater depth helps capture more complex patterns. Furthermore, increasing entries per layer from 256 to 1024 and the code dimension from 128 to 256—for example, moving from ”2share+2specific” (256x128) to ”3share+
2specific” (1024x256)—produces substantial additional improvements.
| Model | Data Volume | Training Time | Storage |
| FOUND | 20,000,000 samples | about 7 hours | 55GB |
| U2QT | 20,000,000 samples | about 2 hours | 0.65GB |
5.4. Efficiency Analysis
In addition to predictive performance, we compare the computational efficiency of U²QT and the FOUND baseline in terms of training time and storage (Table 6), using a dataset of 20 million samples. U²QT trains 3.5 times faster (2 hours vs. 7 hours) and produces user representations that are over 84 times smaller (0.65GB vs. 55GB). These results demonstrate that U²QT is not only effective but also highly efficient and scalable, making it well-suited for large-scale, real-world applications where computational and storage resources are limited.
5.5. Codebook Analysis
5.5.1. CodeBook Utilization
In this section, we examine the utilization rates of the shared and domain-specific codebooks across the two quantization layers to better understand our model’s behavior. We define the utilization rate as the percentage of unique codes used during training, which provides insight into how effectively the model leverages its representational capacity (van der Maaten and Hinton, 2008). The results are visualized in Figure 5. From this analysis, we draw two key observations:
i) Hierarchical abstraction and shared-specific disentanglement. The observed utilization patterns indicate that our hierarchical, modular design organizes knowledge across levels of abstraction. The lower utilization of the shared codebook—especially at the first layer—suggests that a small set of cross-domain patterns is captured and separated from domain-specific ones. In contrast, the consistently higher utilization of the second layer and the domain-specific codebooks reflects the model’s ability to refine and recombine fundamental patterns into diverse, fine-grained domain representations. This synergy between hierarchical abstraction and knowledge separation supports the design’s effectiveness in representing both broad and nuanced information efficiently.
ii) Domain-dependent capacity usage. Differences in utilization across domains align with the inherent properties of the data. For example, the ”Tabular” domain exhibits the lowest utilization, which is expected because tabular data, composed primarily of numerical fields, tends to have less semantic richness and diversity than textual data from domains such as ”App” or ”Bill.” As a result, it can be represented well by a smaller, more specialized set of codes, and the observed utilization patterns faithfully reflect these underlying characteristics.
5.5.2. Multi-source Visualization
Figure 6 presents a T-SNE visualization comparing the domain embeddings learned by our U²QT model (b), which employs a 2 shared + 2 specific codebook architecture, against a baseline variant using only four shared codebook layers (a). The comparison reveals a stark contrast. In Figure 6(b), our proposed architecture produces distinct, well-separated clusters for each domain. Each cluster exhibits high intra-domain cohesion (tightly packed points) and clear inter-domain separability (significant distance between clusters). Conversely, the purely shared architecture shown in Figure 6(a) struggles to achieve this separation. The resulting clusters are visibly more dispersed and less defined, with ambiguous boundaries between several domains (e.g., Search and SPM). This qualitative evidence strongly validates our shared+specific design, proving its necessity for learning high-quality representations from heterogeneous data sources.
6. Conclusion
In this paper, we propose Unified User Tokenizer (U2QT), a novel framework for user representation. To unify the representation of heterogeneous user behaviors from various sources, we utilize Qwen3 embedding model to encode them in language space. For more efficient storage and computation, we design a multi-view RQ-VAE that compresses user representations into discrete tokens. We conduct comprehensive experiments across various downstream tasks to evaluate the effectiveness of our approach. The results consistently demonstrate that U2QT delivers strong representation capability, efficient storage and computation, and flexible cross-task generalization.
Appendix A Baselines
We compare our U2QT with the following state-of-the-art methods in user representation.
MSDP (Fu et al., 2023): Multi-scale Stochastic Distribution Prediction model for learning user behavioral sequence representation, which takes the prediction on user’s behaviors distribution over a period of time as the self-supervision signal.
One4all (Shin et al., 2021):One4all is a general-purpose representation learning through large-scale pre-training, while U-MLP is one of its extended user targeting model which adds MLP user decoder to generate targeted users.
FOUND (Dou et al., 2025):FOUND is a general user modeling method pretrained by self-supervised stage and user-text alignment stage, which enables zero-shot transfer and few-shot user targeting.
MSDP+Unified Tokenizer: We take place the raw user behavior data with our pretrained unified user tokenizers, which is able to verify the necessity of early-fusion strategy in multi-source fusion.
One4all+Unified Tokenizer: The same operations and purpose as above.
We also utilize the pretrained user tokenizer in downstream recommendation task based on following model,
MaskNet (Wang et al., 2021): MaskNet introduces multiplicative operation into DNN ranking system by proposing instance-guided mask which performs element-wise product both on the feature embedding and feedforward layers.
Appendix B Evaluation Metrics
For User Representation including future behavior and user profile prediction, linear probe representation analysis is conducted on all annotated datasets, which is evaluated by AUC (Area Under the ROC Curve (Bradley, 1997)) and KS (Kolmogorov-Smirnov (Berger and Zhou, 2014)) metrics. For Recommendation Task, we utilize AUC and HR (Hit Rate) to evaluate the performance of models.
Appendix C Implementation Detail
In this section, we present the detailed implementation of our U2QT framework. We first describe the hyperparameters used in our model. First, we utilize the Qwen3-Embedding-0.6B222https://huggingface.co/Qwen/Qwen3-Embedding-0.6B to derive embeddings in 1024-dim for mutli-source data. The multi-view RQ-VAE module consists of a two-layer shared codebook with 256 entries of dimension 768, as well as six two-layer source-specific codebooks comprising 256 entries of dimension 128. The model is optimized using AdamW with a learning rate of 1e-3. All experiments are conducted on a cluster of sixteen PPU 810e processors, each with 100 GB of memory.